OpenStudy (anonymous):

What is the side length of a cube with volume 4913 cm3?

OpenStudy (anonymous):

x^3 = 4913 Raise both sides to the (1/3) power x = 4913 ^ (1/3) Put that in your calculator...or if you have a cube root key, take the cube root of 4913